Problem 23.2A Computing financial ratios. LO 23-5, 23-6, 23-7 The Stair Company sells computer parts through a retall sore that it operates. The firms comparative income statement and balance sheet for the years 2019 and 2018 are shown belovw THE STAIR COMPANY Comparabire Income 3sutement For Years End-d Decembex 01, 2019 and ao18 Amount 2019201 Revenue Bales Les, 3.1es Return and Al2owance , 1,127,500 ธ18,400 27.500 18,400 1,100,000 00,000 Net ales Cost of Goods 3old Merchandise Inventory, January 1 Net Parchase Total Mexehandine Availabie fox Sale Lese Merahandise Invencory, December a1 110,000 soo, eod 348,150 254400 458,150 355,200 115,500 110,000 a eso 245 200 787. as0 858.800 Cost of Goods 3old Gross Profis on Sales Operasing Expenses 3 11ing Expenses 99.000 80,000 8 8008.000 a1: 900 20,800 129 700 200-800 Payxoll Tas Expense-Jelling Othes 3Selling Expenses Total 3e1ling Expenses Genexal and Administrative Expenses Officiers, 3교iarien Expense 157,300 132, 000 15, 400 9,600 8, 8008,800 14. 300 9,600 95 800 140,000 a35,500 348800 421. 850 a06.000 16555 91.600 295,295 214,200 Depreeiation Expense Other Genexal and Adminintzavive Expenses Total Genezal and Adminintrabive Expenses Totsl Operating Expenses les Income befoze Income Taxes Income Tax Expense det Income aftex Income Taxes
media%2F1be%2F1beeeb34-6ad0-4394-9824-68
media%2F40b%2F40b068ac-d0fe-4f39-a668-3f
Answer #1
  1. Current ratio: current assets/ current liabilities

2018

2019

CA=276200/87100

=3.17

CA= 362300/49700

=7.28

  1. Acid-test ratio =(current assets-inventories)/current liabilities

2018

2019

=(276200-110000) /87100

=1.90

=(362300-115500)/49700

=4.96

  1. Inventory turnover ratio= cost of goods sold/average inventory

Average inventory= (opening inventory + closing inventory)/2

2018 average inventory= (100800+110000)/2=105400

2019 average inventory =(110000+115500)/2=112750

2018

2019

=245200/105400=2.32times

=242650/112750=2.15 times

  1. Return on sales= net income(after tax)/net sales

2018

2019

=214200/800000=.267

=295295/1100000 =.268

  1. Earning per share =net income/no . of common stock

2018

2019

=245200/20000 =12.26

=295295/20000=14.76

  1. Book value per share = total common stock holders equity/no of common shares

2018

2019

=20000/20000=1

=20000/20000 =1

  1. Return on total asset= EBIT(earning before interest and tax)/total asset

EBIT 2018=206000 EBIT 2019=421650

2018

2019

=206000/435000=.47

=421650/512300=.82